determine the length of the sides of a cube who's interior DIAGONALS measure 173mm.
Answer by checkley77(12569)

A right triangle with one leg=x, the other leg=xsqrt2 & a hypotenuse=173
x^2+(xsqrt2)^2=173
x^2+2x^2=173
3x^2=173
x^2=173/3
x^2=57.6667
x=sqrt57.6667
x=7.594 ans. for the diagonal.
Proof:
7.594^2+7.594^2*2=173
57.6667+57.6667*2=173
57.6667+115.333=173
173=173